# Asymptotic directions in the moduli space of curves

### v1

@@ 9. Rank 2 deformations
In fact, if $h^{0}(L^{\otimes 2})>4$, $C$ has a $g^{3}_{8}$. This can’t give a birational map, by the Castelnuovo bound. So the map has degree 2 and it gives a double cover of genus 1 curve, which contradicts our assumptions. Hence $h^{0}(L^{\otimes 2})=3$.
If $h^{0}(L^{\otimes 3})>4$, $C$ has a $g^{4}_{12}$. Again the induced map can’t be birational by the Castelnuovo bound. If it has degree 3, the curve $C$ would be trigonal, which contradicts our assumption on $Cliff(C)$. If the map has degree 2, it gives a double cover of a genus 1 or 2 curve, a contradiction.
@@
Assume the curve $C$ is bielliptic of genus at least 5, and let $\pi:C\to E$ be the $2:1$ map to a genus 1 curve $E$. By the Castelnuovo inequality a bielliptic curve of genus at least 5 is not hyperelliptic, nor trigonal. Denote by $\sigma$ the bielliptic involution. Consider the curve $\Gamma$ in the surface $S=C\times C$ given by the graph of $\sigma$, $\Gamma:=\{(p,\sigma(p))\ |\ p\in C\}$. Consider the form $\hat{\eta}\in H^{0}(K_{S}(2\Delta))$ introduced in [3]. Denote by $Z(\hat{\eta})$ its zero locus. Clearly $Z(\hat{\eta})$ intersects the curve $\Gamma$ outside the diagonal. Take a point $(p,\sigma(p))\in Z(\hat{\eta})\cap\Gamma$, $p\neq\sigma(p)$. Thus, for any quadric $Q\in I_{2}$, by [3] we have
@@
Notice that any quadric $Q\in I_{2}$, $Q$ is $\sigma$-invariant. In fact, $H^{0}(C,K_{C})\cong H^{0}(E,K_{E})\oplus H^{0}(C,K_{C})^{-}$, where $H^{0}(C,K_{C})^{-}$ denotes the anti-invariant subspace by the action of $\sigma$ and $H^{0}(E,K_{E})\cong H^{0}(C,K_{C})^{+}$ is the invariant subspace. Hence $\operatorname{dim}\big{(}Sym^{2}H^{0}(C,K_{C})\big{)}^{-}=g-1$. Moreover we have $\operatorname{dim}H^{0}(C,K_{C}^{\otimes 2})^{+}=2g-2$, thus $\operatorname{dim}H^{0}(C,K_{C}^{\otimes 2})^{-}=g-1$. Since the multiplication map is $\sigma$-equivariant, $I_{2}(K_{C})^{-}=(0)$, so $I_{2}(K_{C})=I_{2}(K_{C})^{+}$. Then we have
@@
The asymptotic directions found in Theorem 9.2 are linear combination of Schiffer variations supported on a divisor $D=2p+2q$ of a $g^{1}_{4}$ on the bielliptic curve $C$.
@@
Then by a Theorem of Macaulay (see [14, Theorem 2]), we have $c_{1}\leq 2$. Gotzmann’s Persistence Theorem (see [14]) says that if $c_{1}=2$, then $c_{i}=2$ for all $i\geq 1$. From this one easily proves that ${\mathcal{O}}_{Z}$ has length 2.
@@
By Proposition 9.4 we have shown that a rank 2 deformations $\zeta$ that is not a linear combinations of Schiffer variations corresponds to a point $Z\in Hilb^{2}({\mathbb{P}}^{2})$ whose support does not intersect $C$. We have $ker(\zeta)\cong H^{0}({\mathcal{O}}_{{\mathbb{P}}^{2}}(3)\otimes{\mathcal{I}}_{Z})$ and $\zeta$ is determined by its kernel.
The deformation $\zeta$ corresponds to such a length 2 scheme $Z$, that is the intersection of a line $l_{1}$ and a conic $\Gamma$, which are tangent in a point if $Z$ is not reduced. So $ker(\zeta)\cong H^{0}({\mathcal{O}}_{{\mathbb{P}}^{2}}(3)\otimes{\mathcal{I}}_{Z})=l\cdot H^{0}({\mathcal{O}}_{{\mathbb{P}}^{2}}(2))+\Gamma\cdot H^{0}({\mathcal{O}}_{{\mathbb{P}}^{2}}(1))$. The scheme $Z$ is either supported in two distinct points $p$, $q$, or it is given by $p$ and a tangent direction $v$. So the line $l_{1}$ is uniquely determined by $Z$, while we can choose $\Gamma$ as the union of two lines $l_{2},l_{3}$ passing through $p$ and $q$, or passing through $p$ if $Z$ is not reduced. Denote by $l_{i{|C}}=:s_{i}\in H^{0}({\mathcal{O}}_{C}(1))$, set $\omega=s_{1}s_{2}s_{3}\in H^{0}(K_{C})$ and $D:=div(s_{1})$. Then a Dolbeault representative for $\zeta$ is ¯$\overline{\partial}(\frac{\rho_{D}}{\omega})$, where $\rho_{D}$ is as in Proposition 2.3. This can be easily seen, observing that $ker([\overline{\partial}(\frac{\rho_{D}}{\omega})])$ is equal to $ker(\zeta)$. So $\zeta$ is double split, hence it is not asymptotic by Proposition 6.2.
@@
Consider $s_{1}\in H^{0}({\mathcal{O}}_{C}(1)(-p-q))$, $s_{2}\in H^{0}({\mathcal{O}}_{C}(1))$ such that $\langle s_{1},s_{2}\rangle$ is base point free and such that $p$ is not a ramification point of this pencil. Let $s_{3}\in H^{0}({\mathcal{O}}_{C}(1))$, with $s_{3}(q)=0$, $s_{3}(p)\neq 0$. Set $Q=s_{1}^{2}s_{3}\odot s_{2}^{2}s_{3}-s_{1}s_{2}s_{3}\odot s_{1}s_{2}s_{3}\in I_{2}(K_{C})$. Then by [2, Lemma 2.2] we have
@@
so $\mu_{2}(Q)(p)\neq 0$, while $\mu_{2}(Q)(q)=0$ and $Q(p,q)=0$. So by [3, Thm. 2.2], we have $II(Q)(\zeta)=\pi ia^{2}\mu_{2}(Q)(p)=0$ if and only if $a=0$ and $\zeta$ is a Schiffer variation at $q$, so it is not asymptotic.
If the line through $p$ and $q$ is bitangent, take $s_{2}$ the section of $H^{0}({\mathcal{O}}_{C}(1))$ given by a line passing through $p$ and not through $q$, $s_{3},s_{4}$ the sections given by distinct lines passing through $q$ and not through $p$. Then the quadric $Q=s_{2}^{2}s_{4}\odot s_{3}^{2}s_{4}-s_{2}s_{3}s_{4}\odot s_{2}s_{3}s_{4}\in I_{2}(K_{C})$, is such that $Q(p,q)=0$, since $s_{2}s_{4}(p)=s_{2}s_{4}(q)=0$. By [2, Lemma 2.2], we have
@@
Hence $\mu_{2}(Q)(p)\neq 0$, while $\mu_{2}(Q)(q)=0$. So by [3, Thm. 2.2], we have $II(Q)(\zeta\odot\zeta)=-2\pi ia^{2}\mu_{2}(Q)(p)=0$ if and only if $a=0$ and $\zeta$ is a Schiffer variation at $q$, so it is not asymptotic.
@@ 10. Maroni special trigonal curves of genus $6,7$ and trigonal curves of genus $5$
The space $W$ of quadrics that contain the canonical curve is spanned by the rank $\leq 4$ quadrics corresponding to all the pencils of $H^{0}(M)$. Using the chosen basis and computing the second fundamental form as in Proposition 4.3, we easily see $II(Q)(\zeta\odot\zeta)=0$, for all the quadrics constructed as above, except for (possibly) the ones corresponding to the three pencils in the subspace $\langle\sigma s_{2}^{2},\sigma s_{1}s_{2},t\rangle\subset H^{0}(M)$. Denote by
@@
Now, by Riemann Roch, $h^{0}(M^{\otimes 3}\otimes L)=24-5=19.$ Note that $M^{\otimes 3}\otimes L\cong L^{\otimes 7}(3q).$ Set $V=H^{0}(L)$, then $s_{1}^{i}\cdot s_{2}^{n-i}$ is a basis of $Sym^{n}(V)$ the space of symmetric tensor of $V.$ Consider the space
@@
Since the curve is trigonal (non hyperelliptic) $\phi_{1}(s_{1},s_{2})\neq 0$ and since $\sigma\phi_{3}(s_{1},s_{2})t^{2}+\sigma^{2}\phi_{5}(s_{1},s_{2})+\sigma^{3}\phi_{7}(s_{1},s_{2})$ vanishes on $q$ we may take $\phi_{1}(s_{1},s_{2})=s_{2}$ then we get the relation:
@@
Setting $t^{\prime}:=t+\alpha\sigma s_{1}^{2}+\beta\sigma s_{1}s_{2}$, we can assume that $\phi_{3}(s_{1},s_{2})=as_{1}^{3}+bs_{2}^{3}$, with $a\neq 0$.
@@
Moreover, since $a\neq 0$, changing $y$ by a non zero multiple, we can assume that $a=1$.
@@
We observe that equation (18) depends on 11 parameters, while the dimension of the locus of trigonal curves of genus 6 with Maroni degree 1 has dimension 12.
By [1, Lemma 3.3], the tangent space to the trigonal locus is:

### v2

@@ 9. Rank 2 deformations
In fact, if $h^{0}(L^{\otimes 2})\geq 4$, $C$ has a $g^{3}_{8}$. This can’t give a birational map, by the Castelnuovo bound. The map cannot have degree $4$ otherwise the image would be a conic, hence contained in a plane. So the map has degree 2 and the image has degree 4, so it is either rational or a genus 1 curve. Since it is not hyperelliptic, nor bielliptic by assumption, we have $h^{0}(L^{\otimes 2})=3$.
If $h^{0}(L^{\otimes 3})\geq 5$, $C$ has a $g^{4}_{12}$. Again the induced map can’t be birational by the Castelnuovo bound. If it has degree 3, since it is non degenerate, the image is a rational normal curve of degree 4, hence the curve $C$ is trigonal, which contradicts our assumption on $Cliff(C)$. If the map has degree 2, the image is non degenerate of degree 6, hence by Clifford’s theorem it is non special, hence by Riemann Roch it is a curve of genus 1 or 2, a contradiction.
@@
Assume the curve $C$ is bielliptic of genus at least 5, and let $\pi:C\to E$ be the $2:1$ map to a genus 1 curve $E$. By the Castelnuovo-Severi inequality (see e.g. [1, Theorem 3.5]), a bielliptic curve of genus at least 5 is not hyperelliptic, nor trigonal.
Denote by $\sigma$ the bielliptic involution. Consider the curve $\Gamma$ in the surface $S=C\times C$ given by the graph of $\sigma$, $\Gamma:=\{(p,\sigma(p))\ |\ p\in C\}$. Consider the form $\hat{\eta}\in H^{0}(K_{S}(2\Delta))$ introduced in [7]. Denote by $Z(\hat{\eta})$ its zero locus. Clearly $Z(\hat{\eta})$ intersects the curve $\Gamma$ outside the diagonal. Take a point $(p,\sigma(p))\in Z(\hat{\eta})\cap\Gamma$, $p\neq\sigma(p)$. Thus, for any quadric $Q\in I_{2}$, by [7] we have
@@
Notice that any quadric $Q\in I_{2}$ is $\sigma$-invariant. In fact, $H^{0}(C,K_{C})\cong H^{0}(E,K_{E})\oplus H^{0}(C,K_{C})^{-}$, where $H^{0}(C,K_{C})^{-}$ denotes the anti-invariant subspace by the action of $\sigma$ and $H^{0}(E,K_{E})\cong H^{0}(C,K_{C})^{+}$ is the invariant subspace. Hence $\operatorname{dim}\big{(}Sym^{2}H^{0}(C,K_{C})\big{)}^{-}=g-1$. Moreover we have $\operatorname{dim}H^{0}(C,K_{C}^{\otimes 2})^{+}=2g-2$, thus $\operatorname{dim}H^{0}(C,K_{C}^{\otimes 2})^{-}=g-1$. Since the multiplication map is $\sigma$-equivariant, $I_{2}(K_{C})^{-}=(0)$, so $I_{2}(K_{C})=I_{2}(K_{C})^{+}$. Then we have
@@
The asymptotic directions found in Theorem 9.2 are linear combinations of Schiffer variations supported on a divisor $D=2p+2q$ of a $g^{1}_{4}$ on the bielliptic curve $C$.
@@
Then by a Theorem of Macaulay (see [18, Theorem 2]), we have $c_{1}\leq 2$. Gotzmann’s Persistence Theorem (see [18]) says that if $c_{1}=2$, then $c_{i}=2$ for all $i\geq 1$. From this one easily proves that ${\mathcal{O}}_{Z}$ has length 2.
@@
By Proposition 9.4 we have shown that a rank 2 deformation $\zeta$ that is not a linear combination of Schiffer variations corresponds to a point $Z\in Hilb^{2}({\mathbb{P}}^{2})$ whose support does not intersect $C$. We have $ker(\zeta)\cong H^{0}({\mathcal{O}}_{{\mathbb{P}}^{2}}(3)\otimes{\mathcal{I}}_{Z})$ and $\zeta$ is determined by its kernel.
The deformation $\zeta$ corresponds to such a length 2 scheme $Z$, that is the intersection of a line $l_{1}$ and a conic $\Gamma$, which are tangent in a point if $Z$ is not reduced. So $ker(\zeta)\cong H^{0}({\mathcal{O}}_{{\mathbb{P}}^{2}}(3)\otimes{\mathcal{I}}_{Z})=l\cdot H^{0}({\mathcal{O}}_{{\mathbb{P}}^{2}}(2))+\Gamma\cdot H^{0}({\mathcal{O}}_{{\mathbb{P}}^{2}}(1))$. The scheme $Z$ is either supported in two distinct points $p$, $q$, or it is given by $p$ and a tangent direction $v$. So the line $l_{1}$ is uniquely determined by $Z$, while we can choose $\Gamma$ as the union of two lines $l_{2},l_{3}$ passing through $p$ and $q$, or passing through $p$ if $Z$ is not reduced. Denote by $l_{i{|C}}=:s_{i}\in H^{0}({\mathcal{O}}_{C}(1))$, set $\omega=s_{1}s_{2}s_{3}\in H^{0}(K_{C})$ and $D:=div(s_{1})$. Then a Dolbeault representative for $\zeta$ is ¯$\overline{\partial}(\frac{\rho_{D}}{\omega})$, where $\rho_{D}$ is as in Proposition 2.3. This can be easily seen, observing that $ker([\overline{\partial}(\frac{\rho_{D}}{\omega})])$ is equal to $ker(\zeta)$. In fact, $ker([\overline{\partial}(\frac{\rho_{D}}{\omega})])$ is equal to $s_{1}\cdot H^{0}({\mathcal{O}}_{C}(2))+s_{1}s_{2}\cdot H^{0}({\mathcal{O}}_{C}(1))$, as it is easily seen using Proposition 2.3.
So $\zeta$ is double split, hence it is not asymptotic by Proposition 6.3.
@@
Consider $s_{1}\in H^{0}({\mathcal{O}}_{C}(1)(-p-q))$, $s_{2}\in H^{0}({\mathcal{O}}_{C}(1))$ such that $\langle s_{1},s_{2}\rangle$ is base point free and such that $p$ is not a ramification point of this pencil. Let $s_{3}\in H^{0}({\mathcal{O}}_{C}(1))$, with $s_{3}(q)=0$, $s_{3}(p)\neq 0$. Set $Q=s_{1}^{2}s_{3}\odot s_{2}^{2}s_{3}-s_{1}s_{2}s_{3}\odot s_{1}s_{2}s_{3}\in I_{2}(K_{C})$. Then by [6, Lemma 2.2] we have
@@
so $\mu_{2}(Q)(p)\neq 0$, while $\mu_{2}(Q)(q)=0$ and $Q(p,q)=0$. So by [7, Thm. 2.2], we have $II(Q)(\zeta\odot\zeta)=\pi ia^{2}\mu_{2}(Q)(p)=0$ if and only if $a=0$ and $\zeta$ is a Schiffer variation at $q$, so it is not asymptotic.
If the line through $p$ and $q$ is bitangent, take $s_{2}$ the section of $H^{0}({\mathcal{O}}_{C}(1))$ given by a line passing through $p$ and not through $q$, $s_{3},s_{4}$ the sections given by distinct lines passing through $q$ and not through $p$. Then the quadric $Q=s_{2}^{2}s_{4}\odot s_{3}^{2}s_{4}-s_{2}s_{3}s_{4}\odot s_{2}s_{3}s_{4}\in I_{2}(K_{C})$, is such that $Q(p,q)=0$, since $s_{2}s_{4}(p)=s_{2}s_{4}(q)=0$. By [6, Lemma 2.2], we have
@@
Hence $\mu_{2}(Q)(p)\neq 0$, while $\mu_{2}(Q)(q)=0$. So by [7, Thm. 2.2], we have $II(Q)(\zeta\odot\zeta)=-2\pi ia^{2}\mu_{2}(Q)(p)=0$ if and only if $a=0$ and $\zeta$ is a Schiffer variation at $q$, so it is not asymptotic.
@@ 10. Maroni special trigonal curves of genus $6,7$ and trigonal curves of genus $5$
The space $W$ of quadrics that contain the canonical curve is spanned by the rank $\leq 4$ quadrics corresponding to all the pencils of $H^{0}(M)$ (see Remark 6.1). Using the chosen basis and computing the second fundamental form as in Proposition 4.4, we easily see $II(Q)(\zeta\odot\zeta)=0$, for all the quadrics constructed as above, except for (possibly) the ones corresponding to the three pencils in the subspace $\langle\sigma s_{2}^{2},\sigma s_{1}s_{2},t\rangle\subset H^{0}(M)$. Denote by
@@
Now, by Riemann Roch, $h^{0}(M^{\otimes 3}\otimes L)=24-5=19.$ Note that $M^{\otimes 3}\otimes L\cong L^{\otimes 7}(3q).$ Set $V=H^{0}(L)$, then $s_{1}^{i}\cdot s_{2}^{n-i}$ is a basis of $Sym^{n}(V)$ the space of symmetric tensors of $V.$ Consider the space
@@
Since the curve is trigonal (non hyperelliptic) $\phi_{1}(s_{1},s_{2})\neq 0$ and since $\sigma\phi_{3}(s_{1},s_{2})t^{2}+\sigma^{2}\phi_{5}(s_{1},s_{2})t+\sigma^{3}\phi_{7}(s_{1},s_{2})$ vanishes on $q$ we may take $\phi_{1}(s_{1},s_{2})=s_{2}$ then we get the relation:
@@
By a suitable change of coordinates of the form $t^{\prime}:=t+\alpha\sigma s_{1}^{2}+\beta\sigma s_{1}s_{2}$, we can assume that $\phi_{3}(s_{1},s_{2})=as_{1}^{3}+bs_{2}^{3}$, with $a\neq 0$.
@@
Moreover, since $a\neq 0$, changing $t$ (hence $y$) by a non zero multiple, we can assume that $a=1$.
@@
Since in equation (19), we have $\psi_{7}(0)\neq 0$, changing $s_{1}$ and $s_{2}$ (hence $x$ and $y$) by suitable non-zero multiples, we can assume $\psi_{7}(0)=1$. Hence equation (19) depends on 11 parameters, while the dimension of the locus of trigonal curves of genus 6 with Maroni degree 1 has dimension 12.
By [5, Lemma 3.3], the tangent space to the trigonal locus is: